# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ Appendix A Proofs
Case 2: Suppose $|X|\geq|Z|\geq|Y|$. Let $\pi_{1}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(X,Z)$
$\pi_{1}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{k}||\mathbf{v}_{\pi(i)}-\mathbf{z}_{i}||+\sum_{i=k+1}^{m}||\mathbf{v}_{\pi(i)}||\bigg{]}$
Likewise, let $\pi_{2}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(Z,Y)$
$\pi_{2}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{k}}\bigg{[}\sum_{i=1}^{n}||\mathbf{z}_{\pi(i)}-\mathbf{u}_{i}||+\sum_{i=n+1}^{k}||\mathbf{z}_{\pi(i)}||\bigg{]}$
Then, we have
$d_{M}(X,Z)+d_{M}(Z,Y)$ $=\sum_{i=1}^{k}||\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{z}_{i}||+\sum_{i=k+1}^{m}||\mathbf{v}_{\pi_{1}^{*}(i)}||+\sum_{i=1}^{n}||\mathbf{z}_{\pi_{2}^{*}(i)}-\mathbf{u}_{i}||+\sum_{i=n+1}^{k}||\mathbf{z}_{\pi_{2}^{*}(i)}||$
$=\sum_{i=1}^{n}\Big{[}||\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}-\mathbf{z}_{\pi_{2}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(i)}-\mathbf{u}_{i}||\Big{]}$
$\qquad\qquad\qquad+\sum_{i=n+1}^{k}\Big{[}||\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}-\mathbf{z}_{\pi_{2}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(i)}||\Big{]}+\sum_{i=k+1}^{m}||\mathbf{v}_{\pi_{1}^{*}(i)}||$
$\geq\sum_{i=1}^{n}||\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}-\mathbf{u}_{i}||+\sum_{i=n+1}^{k}||\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}||+\sum_{i=k+1}^{m}||\mathbf{v}_{\pi_{1}^{*}(i)}||$
$\geq\min_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi(i)}||\bigg{]}$
@@
Case 3: Suppose $|Z|\geq|Y|\geq|X|$. Let $\pi_{1}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(X,Z)$
$\pi_{1}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{k}}\sum_{i=1}^{m}\bigg{[}||\mathbf{v}_{i}-\mathbf{z}_{\pi(i)}||+\sum_{i=m+1}^{k}||\mathbf{z}_{\pi(i)}||\bigg{]}$
Likewise, let $\pi_{2}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(Z,Y)$
$\pi_{2}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{k}}\bigg{[}\sum_{i=1}^{n}||\mathbf{u}_{i}-\mathbf{z}_{\pi(i)}||+\sum_{i=n+1}^{k}||\mathbf{z}_{\pi(i)}||\bigg{]}$
Then, we have
$d_{M}(X,Z)+d_{M}(Z,Y)=\sum_{i=1}^{m}||\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}||+\sum_{i=m+1}^{k}||\mathbf{z}_{\pi_{1}^{*}(i)}||+\sum_{i=1}^{n}||\mathbf{z}_{\pi_{2}^{*}(i)}-\mathbf{u}_{i}||+\sum_{i=n+1}^{k}||\mathbf{z}_{\pi_{2}^{*}(i)}||$
For each $i\in[k]$, there exists a single $j\in[k]$ such that $\pi_{1}^{*}(i)=\pi_{2}^{*}(j)$. For each $i,j\in[k]$ with $\pi_{1}^{*}(i)=\pi_{2}^{*}(j)$ one of the following holds
@@
    $||\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(j)}-\mathbf{u}_{j}||\geq||\mathbf{v}_{i}-\mathbf{u}_{j}||$ if $i\leq m$ and $j\leq n$
@@
    $||\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(j)}||\geq||\mathbf{v}_{i}||$ if $i\leq m$ and $j>n$
@@
    $||\mathbf{z}_{\pi_{1}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(j)}-\mathbf{u}_{j}||\geq||\mathbf{u}_{j}||$ if $i>m$ and $j\leq n$
@@
    $||\mathbf{z}_{\pi_{1}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(j)}||\geq 0$ if $i>m$ and $j>n$
Note that $d_{M}(X,Z)+d_{M}(Z,Y)$ can be written as a sum of $k$ terms, where each term corresponds to one of the above $4$ sums of norms. If we take pairs of terms of types $2$ and $3$ and we sum them, we have that
$||\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(j)}||+||\mathbf{z}_{\pi_{1}^{*}(i)}||+||\mathbf{z}_{\pi_{2}^{*}(j)}-\mathbf{u}_{j}||\geq||\mathbf{v}_{i}||+||\mathbf{u}_{j}||=||\mathbf{v}_{i}||+||-\mathbf{u}_{j}||\geq||\mathbf{v}_{i}-\mathbf{u}_{j}||$
Note also that type $2$ occurs $m-n$ times more than type $3$. Therefore, using the inequalities for the $4$ types of sums of norms above, we have
$d_{M}(X,Z)+d_{M}(Z,Y)$ $\geq\min_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{\pi(i)}||+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi(i)}||\bigg{]}$
@@
A.2 Proof of Proposition 2.3
If $|X|=|Y|=M$, the second and third constraints of the oprimization problem that needs to be solved to compute $d_{\text{EMD}}(X,Y)$ become as follows:
$\begin{split}&\sum_{j=1}^{M}[\mathbf{F}]_{ij}=\frac{1}{M},\quad 1\leq i\leq M\qquad\text{and}\qquad\sum_{i=1}^{M}[\mathbf{F}]_{ij}=\frac{1}{M},\quad 1\leq j\leq M\\ \end{split}$
Therefore, matrix $\mathbf{F}$ is a doubly stochastic matrix. The Birkhoff-von Neumann Theorem states that the set of $M\times M$ doubly stochastic matrices forms a convex polytope whose vertices are the $M\times M$ permutation matrices. Furthermore, it is known that the optimal value of a linear objective in a nonempty polytope is attained at a vertex of the polytope [4]. The optimal solution would thus be a permutation matrix $\mathbf{F}$. Let also $\sigma\in S_{M}$ denote the permutation that is associated with that matrix. Therefore, we have that:
$\begin{split}Md_{\text{EMD}}(X,Y)&=M\min_{\mathbf{P}\in\mathcal{B}_{M}}\sum_{i=1}^{M}\sum_{j=1}^{M}[\mathbf{P}]_{ij}\,||\mathbf{v}_{i}-\mathbf{u}_{j}||_{2}\\ &=M\min_{\mathbf{P}\in\boldsymbol{\Pi}_{M}}\sum_{i=1}^{M}\sum_{j=1}^{M}[\mathbf{P}]_{ij}\,||\mathbf{v}_{i}-\mathbf{u}_{j}||_{2}\\ &=M\min_{\pi\in\mathfrak{S}_{M}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||_{2}\bigg{]}\\ &=d_{M}(X,Y)\end{split}$
A.3 Proof of Theorem 3.1
A.3.1 The mean function is Lipschitz continuous with respect to EMD
Let $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}\!\!\}$ and $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}\!\!\}$ be two multisets, consisting of $m$ and $n$ vectors of dimension $d$, respectively. Then, we have that:
$\Big{|}\Big{|}f_{\text{mean}}(X)-f_{\text{mean}}(Y)\Big{|}\Big{|}$ $=\bigg{|}\bigg{|}\frac{1}{m}\sum_{i=1}^{m}\mathbf{v}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\sum_{i=1}^{m}\frac{1}{m}\mathbf{v}_{i}-\sum_{j=1}^{n}\frac{1}{n}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\sum_{i=1}^{m}\bigg{(}\sum_{j=1}^{n}[\mathbf{F}]_{ij}\bigg{)}\mathbf{v}_{i}-\sum_{j=1}^{n}\bigg{(}\sum_{i=1}^{m}[\mathbf{F}]_{ij}\bigg{)}\mathbf{u}_{j}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}]_{ij}(\mathbf{v}_{i}-\mathbf{u}_{j})\bigg{|}\bigg{|}$
$\leq\sum_{i=1}^{m}\sum_{j=1}^{n}||[\mathbf{F}]_{ij}(\mathbf{v}_{i}-\mathbf{u}_{j})||$
$=\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}]_{ij}||(\mathbf{v}_{i}-\mathbf{u}_{j})||$
@@ A.3.1 The mean function is Lipschitz continuous with respect to EMD
A.3.2 The mean function is not Lipschitz continuous with respect to the matching distance
Suppose that the mean function is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>(2L+1)\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we have $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=||\mathbf{v}_{2}||=\sqrt{d}\epsilon$. We also have that:
$\Big{|}\Big{|}f_{\text{mean}}(X)-f_{\text{mean}}(Y)\Big{|}\Big{|}$ $=\bigg{|}\bigg{|}\frac{1}{2}\sum_{i=1}^{2}\mathbf{v}_{i}-\mathbf{u}_{1}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\frac{1}{2}\mathbf{v}_{1}+\frac{1}{2}\mathbf{v}_{2}-\mathbf{u}_{1}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\frac{1}{2}(c,c,\ldots,c)^{\top}+\frac{1}{2}(\epsilon,\epsilon,\ldots,\epsilon)^{\top}-(c,c,\ldots,c)^{\top}\bigg{|}\bigg{|}$
$=\bigg{|}\bigg{|}\Big{(}\frac{\epsilon-c}{2},\frac{\epsilon-c}{2},\ldots,\frac{\epsilon-c}{2}\Big{)}^{\top}\bigg{|}\bigg{|}$
$=\frac{1}{2}||(\epsilon-c,\epsilon-c,\ldots,\epsilon-c)^{\top}||$

### v3

@@ Appendix B Proofs
Case 2: Suppose $|X|\geq|Z|\geq|Y|$. Let $\pi_{1}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(X,Z)$:
$\pi_{1}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{k}\|\mathbf{v}_{\pi(i)}-\mathbf{z}_{i}\|+\sum_{i=k+1}^{m}\|\mathbf{v}_{\pi(i)}\|\bigg]$
Likewise, let $\pi_{2}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(Z,Y)$:
$\pi_{2}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{k}}\bigg[\sum_{i=1}^{n}\|\mathbf{z}_{\pi(i)}-\mathbf{u}_{i}\|+\sum_{i=n+1}^{k}\|\mathbf{z}_{\pi(i)}\|\bigg]$
Then, we have:
$d_{M}(X,Z)+d_{M}(Z,Y)$ $=\sum_{i=1}^{k}\|\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{z}_{i}\|+\sum_{i=k+1}^{m}\|\mathbf{v}_{\pi_{1}^{*}(i)}\|+\sum_{i=1}^{n}\|\mathbf{z}_{\pi_{2}^{*}(i)}-\mathbf{u}_{i}\|+\sum_{i=n+1}^{k}\|\mathbf{z}_{\pi_{2}^{*}(i)}\|$
$=\sum_{i=1}^{n}\Big[\|\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}-\mathbf{z}_{\pi_{2}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(i)}-\mathbf{u}_{i}\|\Big]$
$\qquad\qquad\qquad+\sum_{i=n+1}^{k}\Big[\|\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}-\mathbf{z}_{\pi_{2}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(i)}\|\Big]+\sum_{i=k+1}^{m}\|\mathbf{v}_{\pi_{1}^{*}(i)}\|$
$\geq\sum_{i=1}^{n}\|\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}-\mathbf{u}_{i}\|+\sum_{i=n+1}^{k}\|\mathbf{v}_{\pi_{1}^{*}(\pi_{2}^{*}(i))}\|+\sum_{i=k+1}^{m}\|\mathbf{v}_{\pi_{1}^{*}(i)}\|$
$\geq\min_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi(i)}\|\bigg]$
@@
Case 3: Suppose $|Z|\geq|Y|\geq|X|$. Let $\pi_{1}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(X,Z)$:
$\pi_{1}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{k}}\sum_{i=1}^{m}\bigg[\|\mathbf{v}_{i}-\mathbf{z}_{\pi(i)}\|+\sum_{i=m+1}^{k}\|\mathbf{z}_{\pi(i)}\|\bigg]$
Likewise, let $\pi_{2}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(Z,Y)$:
$\pi_{2}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{k}}\bigg[\sum_{i=1}^{n}\|\mathbf{u}_{i}-\mathbf{z}_{\pi(i)}\|+\sum_{i=n+1}^{k}\|\mathbf{z}_{\pi(i)}\|\bigg]$
Then, we have:
$d_{M}(X,Z)+d_{M}(Z,Y)=\sum_{i=1}^{m}\|\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}\|+\sum_{i=m+1}^{k}\|\mathbf{z}_{\pi_{1}^{*}(i)}\|+\sum_{i=1}^{n}\|\mathbf{z}_{\pi_{2}^{*}(i)}-\mathbf{u}_{i}\|+\sum_{i=n+1}^{k}\|\mathbf{z}_{\pi_{2}^{*}(i)}\|$
For each $i\in[k]$, there exists a single $j\in[k]$ such that $\pi_{1}^{*}(i)=\pi_{2}^{*}(j)$. For each $i,j\in[k]$ with $\pi_{1}^{*}(i)=\pi_{2}^{*}(j)$ one of the following holds:
@@
    $\|\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(j)}-\mathbf{u}_{j}\|\geq\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$ if $i\leq m$ and $j\leq n$
@@
    $\|\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(j)}\|\geq\|\mathbf{v}_{i}\|$ if $i\leq m$ and $j>n$
@@
    $\|\mathbf{z}_{\pi_{1}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(j)}-\mathbf{u}_{j}\|\geq\|\mathbf{u}_{j}\|$ if $i>m$ and $j\leq n$
@@
    $\|\mathbf{z}_{\pi_{1}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(j)}\|\geq 0$ if $i>m$ and $j>n$
Note that $d_{M}(X,Z)+d_{M}(Z,Y)$ can be written as a sum of $k$ terms, where each term corresponds to one of the above $4$ sums of norms. If we take pairs of terms of types $2$ and $3$ and we sum them, we have that:
$\|\mathbf{v}_{i}-\mathbf{z}_{\pi_{1}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(j)}\|+\|\mathbf{z}_{\pi_{1}^{*}(i)}\|+\|\mathbf{z}_{\pi_{2}^{*}(j)}-\mathbf{u}_{j}\|\geq\|\mathbf{v}_{i}\|+\|\mathbf{u}_{j}\|=\|\mathbf{v}_{i}\|+\|-\mathbf{u}_{j}\|\geq\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$
Note also that type $2$ occurs mnm-n times more than type $3$. Therefore, using the inequalities for the $4$ types of sums of norms above, we have:
$d_{M}(X,Z)+d_{M}(Z,Y)$ $\geq\min_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{\pi(i)}\|+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi(i)}\|\bigg]$
@@
In case $\mathbf{0}$ can be an element of the multisets, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ with $X\neq Y$ such that $d_{M}(X,Y)=0$, i. e., the distance between two distinct points can be equal to 0. The rest of the properties still hold, and thus the matching distance in a pseudometric on $\mathcal{S}(\mathbb{R}^{d})$.
B.2 Proof of Proposition 2.3
If $|X|=|Y|=M$, the second and third constraints of the optimization problem that needs to be solved to compute $d_{\text{EMD}}(X,Y)$ become as follows:
[…]
Therefore, matrix $\mathbf{F}$ is a doubly stochastic matrix. The Birkhoff-von Neumann Theorem states that the set of $M\times M$ doubly stochastic matrices forms a convex polytope whose vertices are the $M\times M$ permutation matrices. Furthermore, it is known that the optimal value of a linear objective in a nonempty polytope is attained at a vertex of the polytope (Bertsimas and Tsitsiklis, 1997). The optimal solution would thus be a permutation matrix $\mathbf{P}\in\bm{\Pi}_{M}$ scaled by $1/M$. Let also $\pi\in\mathfrak{S}_{M}$ denote the permutation that is associated with that matrix. Therefore, we have that:
$\begin{split}Md_{\text{EMD}}(X,Y)&=M\min_{\mathbf{F}\in\mathcal{B}_{M}}\sum_{i=1}^{M}\sum_{j=1}^{M}[\mathbf{F}]_{ij}\,\|\mathbf{v}_{i}-\mathbf{u}_{j}\|_{2}\\ &=M\min_{\mathbf{P}\in\bm{\Pi}_{M}}\sum_{i=1}^{M}\sum_{j=1}^{M}\frac{1}{M}[\mathbf{P}]_{ij}\,\|\mathbf{v}_{i}-\mathbf{u}_{j}\|_{2}\\ &=\min_{\pi\in\mathfrak{S}_{M}}\sum_{i=1}^{M}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|_{2}\\ &=d_{M}(X,Y)\end{split}$
B.3 Proof of Theorem 3.1
B.3.1 The mean function is Lipschitz continuous with respect to EMD
Let $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}\!\!\}$ and $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}\!\!\}$ be two multisets, consisting of $m$ and $n$ vectors of dimension $d$, respectively. Let also $\mathbf{F}^{*}$ denote the matrix that minimizes $d_{\text{EMD}}(X,Y)$. Then, we have that:
$\Big\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\Big\|$ $=\bigg\|\frac{1}{m}\sum_{i=1}^{m}\mathbf{v}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{u}_{j}\bigg\|$
$=\bigg\|\sum_{i=1}^{m}\frac{1}{m}\mathbf{v}_{i}-\sum_{j=1}^{n}\frac{1}{n}\mathbf{u}_{j}\bigg\|$
$=\bigg\|\sum_{i=1}^{m}\bigg(\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}\bigg)\mathbf{v}_{i}-\sum_{j=1}^{n}\bigg(\sum_{i=1}^{m}[\mathbf{F}^{*}]_{ij}\bigg)\mathbf{u}_{j}\bigg\|$
$=\bigg\|\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}(\mathbf{v}_{i}-\mathbf{u}_{j})\bigg\|$
$\leq\sum_{i=1}^{m}\sum_{j=1}^{n}\|[\mathbf{F}^{*}]_{ij}(\mathbf{v}_{i}-\mathbf{u}_{j})\|$
$=\sum_{i=1}^{m}\sum_{j=1}^{n}[\mathbf{F}^{*}]_{ij}\|(\mathbf{v}_{i}-\mathbf{u}_{j})\|$
@@ B.3.1 The mean function is Lipschitz continuous with respect to EMD
B.3.2 The mean function is not Lipschitz continuous with respect to the matching distance
Suppose that the mean function is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>(2L+1)\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}\|=\sqrt{d}\epsilon$. We also have that:
$\Big\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\Big\|$ $=\bigg\|\frac{1}{2}\sum_{i=1}^{2}\mathbf{v}_{i}-\mathbf{u}_{1}\bigg\|$
$=\bigg\|\frac{1}{2}\mathbf{v}_{1}+\frac{1}{2}\mathbf{v}_{2}-\mathbf{u}_{1}\bigg\|$
$=\bigg\|\frac{1}{2}(c,c,\ldots,c)^{\top}+\frac{1}{2}(\epsilon,\epsilon,\ldots,\epsilon)^{\top}-(c,c,\ldots,c)^{\top}\bigg\|$
$=\bigg\|\Big(\frac{\epsilon-c}{2},\frac{\epsilon-c}{2},\ldots,\frac{\epsilon-c}{2}\Big)^{\top}\bigg\|$
$=\frac{1}{2}\|(\epsilon-c,\epsilon-c,\ldots,\epsilon-c)^{\top}\|$